# Constraints on the $\gamma$-parameter for the vacuum solution of Cotton gravity with geodesics and shadows

### v1

@@ III.2 Shadow’s radius and constraints to the $\gamma$ parameter
at $2\sigma$ deviation.
@@
We note that, to quantify the theoretical biases in the calibration factor utilized in the inference of the shadow’s size of Sgr* A, the EHT collaboration performs simulations of black hole images to study a wide range of effects concerning the accretion flow properties and the underlying space-time. In this regard, the collaboration assumes the Kerr metric by default, although they also cover black hole metrics that either deviate from the Kerr space-time or describe alternative geometries, such as wormholes or boson stars. Because their analysis does not include the Cotton geometry, by performing tests with the above constraints we introduce a theoretical bias in our own analysis, since we assume these results to be valid for the CG spacetime. The extent to which the Cotton geometry could modify the above fractional deviation, and the constraints stemming from it, could be evaluated by performing a similar analysis to that by the EHT collaboration. However, such an effort lies well beyond the purpose of this work.
@@
Upon inspection of Fig. 2, we verify that $r_{sh}$ starts to increase around values of $\gamma M\sim 10^{-12}$. Below this value, the CG black hole shadow radius coincides with that of a Schwarzschild black hole. We note that this is qualitatively the same behavior that is evidenced in [43] for the Rindler metric. Actually, the contribution of $\gamma$ to the shadow radius becomes significant when the term $\gamma r\gg 1$, while the upper $2\sigma$ bound in Eq.(27) becomes saturated at a value of approximately $\gamma M\simeq 3.525\times 10^{-12}$. For the values of $\gamma M$ predicted by galaxy radial acceleration rotation curves in [13], or the recent constraints obtained by analyzing CG’s influence on CMB power spectrum and large scale structures [54] (i.e. both estimating $\gamma M\sim 10^{-17}$), the shadow radius of the Cotton black hole lies well within the constraints imposed by Sgr A*’s shadow bounds, though there is no appreciable difference between it and its Schwarzschild counterpart. This result is not surprising, given that $\gamma$ encodes large-distance gravitational effects within the CG framework. We have also found that the Cotton black hole shadow radius remains within the $2\sigma$ bounds for $\gamma M\sim 10^{20}$. Such values describe extremely large and unphysical gravitational acceleration effects, exceeding any other current estimates by several orders of magnitude. For this reason, we have opted for excluding them from the $\gamma$ parameter range in Fig. 2.
@@ IV.1 Exact deflection angle
To proceed with our derivation, it is convenient to make the following variable change in the radial coordinate $u=1/r$ in Eq. (14). This allows us to re-write the photon equation of motion as
@@
$\alpha=\sqrt{\frac{1}{2M}}\left(\int_{u_{R}}^{u_{0}}\frac{du}{\sqrt{G(u)}}+\int_{u_{S}}^{u_{0}}\frac{du}{\sqrt{G(u)}}\right)+\Psi_{R}-\Psi_{S},$ (30)
where $u_{0}=1/r_{0}$ is the inverse of the distance of closest approach, while $u_{R}=1/r_{R}$ and $u_{S}$ are the inverse of the radius of the observer and the source, respectively, and the indices refer to the position of the observer $(R)$ and the source $(S)$. The function $\Psi$ encodes a correction term accounting for the finite distance, and is given by
$\Psi=\arcsin\left(bu\sqrt{A(u)}\right)\,.$ (31)
Note that if we take the limit of $\gamma\rightarrow 0$, and consider also that observer and source are located at infinity we will have
@@
which is the well-known deflection angle formula for the Schwarzschild solution. This change in the definition of the deflection angle is necessary because the Cotton metric is not asymptotically flat. Moving on, the integrals of Eq. (30) can be solved by rewriting the function $G(u)$ as
@@
where $u_{1},u_{2}$ and $u_{3}$ are the roots of the cubic polynomial. We can check that $u_{0}=1/r_{0}$ is one of the roots of this polynomial. We have chosen (for reasons that will become clear later) $u_{2}=1/r_{0}$. Now we can write
$u^{3}-u^{2}\left(u_{1}+u_{3}+\frac{1}{r_{0}}\right)+u\left(u_{1}u_{3}+\frac{u_{1}}{r_{0}}+\frac{u_{3}}{r_{0}}\right)$
$-u_{1}\frac{u_{3}}{r_{0}}=u_{3}-\frac{u^{2}}{2M}-\frac{u\gamma}{2M}+\frac{1}{2Mb^{2}}.$ (34)
@@
$u_{1}+u_{3}+\frac{1}{r_{0}}=\frac{1}{2M},$ (35)
$u_{1}u_{3}+\frac{u_{1}}{r_{0}}+\frac{u_{3}}{r_{0}}=-\frac{\gamma}{2M}.$
Solving the above system we find that $u_{1}$ and $u_{3}$ are given by
@@
respectively. Note that we have two different solutions, one of which is $u_{1}>u_{3}$ and another that $u_{3}>u_{1}$. Since our goal is that these roots fulfill the condition $u_{3}>u_{2}>u_{1}$, we choose the negative sign in the expression $u_{1}$ and the positive sign in the expression $u_{3}$. Now both integrals of Eq. (30) can be put into the form
@@
where $F(\delta,k^{2})$ is an incomplete first-order elliptic integral of the parameters
@@
$\Psi(u)=\sin^{-1}\left(bu\sqrt{1-2Mu+\frac{\gamma}{u}}\right),$ (41)
@@
$\delta(u^{\prime})=\sin^{-1}\Biggl{\{}2\sqrt{2}\Biggl{[}\left(\frac{Mr_{0}(1-u^{\prime}r_{0})\sqrt{r_{0}^{2}c}}{\sqrt{r_{0}^{2}c}+6Mr_{0}-r_{0}^{2}}\right)\times$
×$\times\left(\sqrt{r_{0}^{2}c}-2Mr_{0}(2Rr_{0}+1)+r_{0}^{2}\right)^{-1}\Biggr{]}^{1/2}\Biggr{\}},$ (42)
@@
$r_{0}$ $=$ $\frac{b^{2}\gamma}{3}+\frac{2}{3}b\sqrt{b^{2}\gamma^{2}+3}\times$ (44)
×$\times\cos\left[\frac{1}{3}\cos^{-1}\left(\frac{\left(2b^{4}\gamma^{3}+9b^{2}\gamma-54M\right)}{2b\left(b^{2}\gamma^{2}+3\right)^{3/2}}\right)\right].$
For the sake of the graphical representation of the deflection angle we prefer to use the variable $b^{\prime}=1-b_{c}/b$. This variable is limited to a value between $0\leq b^{\prime}\leq 1$, where zero represents the strong field limit and one the weak field one. In Figs. 3 and 4 we show the deflection angle as a function of the variable $b^{\prime}$ for values of $\gamma$ in the range from $10^{-16}$ to $10^{-9}$. We see that $\alpha$ decreases with increasing $\gamma$ and that there is no significant change in the values of $\alpha$ from $\gamma<10^{-13}$. If we increase the value of the parameter $\gamma$ beyond the order of magnitude of $10^{-9}$, the deflection angle begins to give negative values. Observationally, this value diverges for $b^{\prime}=0$ and tends to zero for $b^{\prime}=1$. We expect our theoretical result to reproduce this behaviour. In other words, for the phenomenon of light deflection, Cotton’s parameter is limited to values smaller than $\gamma=10^{-9}$.
Figure 3: Graphical representation of the deflection angle $\alpha$ as a function of the reduced impact parameter $b^{\prime}$ for values of the CG parameter $\gamma=\{10^{-16},10^{-15},10^{-14},10^{-13}\}$.
Figure 4: Graphical representation of the deflection angle $\alpha$ as a function of the reduced impact parameter $b^{\prime}$ for values of the CG parameter $\gamma=\{10^{-12},10^{-11},10^{-10},10^{-9}\}$.
@@ V Images of black holes with thin accretion disks
Having restricted the $\gamma$-space parameter, we now proceed to generate the images of the Cotton black hole (by pushing the value saturating the upper $2\sigma$ bound of the Sgr A* shadow) and qualitatively compare its appearance to that of a Schwarzschild black hole of the same mass parameter. We consider the setting of a black hole illuminated by an optically (i.e. fully transparent to its own radiation) and geometrically thin accretion disk, when observed at both a face-on inclination (i.e. $\theta_{o}=0^{\circ}$) and at a nearly edge-on inclination (i.e. $\theta_{o}=80^{\circ}$).
We begin our considerations from Eq. (14), which we conveniently re-write in the form
$\frac{d\phi}{dr}=\pm\frac{b}{C(r)}\sqrt{\frac{1}{1-b^{2}V_{\rm eff}}},$ (45)
that describes the variation in a photon’s azimuthal angle as a function of its radial coordinate. By numerically ray-tracing the previous equation for a bunch of trajectories within some range of the impact parameter $b$, we compute the null geodesics from an observer’s screen backwards towards their point of origin. Figure 5 displays the geodesics in the Schwarzschild space-time (left panel) for $M=1$, and in the CG space-time (right panel), the latter with a value of $\gamma M=3.525\times 10^{-12}$. Upon close observation of the figure, one highlights the absence of any significant qualitative differences between the near-horizon lensing properties of the two geometries, under the considered parameter values. In addition to larger distances, the effects of $\gamma$ become more noticeable at much higher orders of magnitude than those considered here, as is discussed in Section IV below.
Figure 5: Null geodesics in the Schwarzschild space-time (left panel) and in the CG space-time as a function of impact parameter $b$ normalized with respect to the mass parameter $M$. The photon trajectories are categorized according to the number of half-turns $n$ around the black hole as the ($n=0$) direct image (black), the $n=1$ photon ring (gold), and the $n=2$ photon ring (red), and with an impact parameter spacing of $1/5$, $1/50$, and $1/500$, respectively. The black circle depicts the inner region to the event horizon. The dashed black circumference represents the photon sphere radius. The black hole equatorial plane is represented by the black vertical line. The observer’s screen is located on the right-side axis.
@@
$I(r;\gamma,\mu,\sigma)=\frac{e^{-\frac{1}{2}\left[\gamma+\text{arcsinh}(\frac{r-\mu}{\sigma})\right]^{2}}}{\sqrt{(r-\mu)^{2}+\sigma^{2}}},$ (46)
@@
For the purpose of generating black hole images we restrict our analysis up to the $n=2$ photon ring, as higher-order photon rings will provide a negligible contribution to the net luminosity (as shall be explained in the next section). We shall suitably adapt the three models introduced in [44] to a spherically symmetric scenario, to simulate the accretion flow around the Schwarzschild and Cotton black holes, with the parameters
GLM1 ::: $\gamma=-\frac{3}{2}\,,\quad\mu=0\,,\qquad\ \sigma=\frac{M}{2}\ ,$ (48)
GLM2 ::: $\gamma=0\,,\qquad\mu=0\,,\qquad\ \sigma=\frac{M}{2}\ ,$ (49)
GLM3 ::: $\gamma=-2\,,\quad\ \mu=\frac{17M}{3}\,,\ \sigma=\frac{M}{4}\ .$ (50)
whose corresponding emission profile is depicted in Fig. 6.
Figure 6: Emission profiles corresponding to the GLM1, GLM2 and GLM3 models described by the choice of parameters in Eqs. (48), (49) and (50), respectively, of the intensity function (46).
@@
Figure 7: Visual appearance, in the impact parameter space, of the Schwarzschild black hole ($1^{st}$ and $3^{rd}$ columns) and the CG black hole ($2^{nd}$ and $4^{th}$ columns), for the GLM3 (top), GLM1 (middle) and GLM2 (bottom) emission profiles, for an observer located at a distance of $r_{o}=1000M$, seen from a face-on orientation and from an inclination of $80^{\circ}$, respectively.
@@ VI Isoradials
In this section, we perform an analysis of isoradial curves in the Cotton and Schwarzschild space-times, to complement the black hole images presented in the previous section. We solve the isoradial equatorial curves corresponding to the $n=0$ emission, following the analytical approximation provided by [58], which states that the arrival point of a light ray may be mapped onto an observer’s plane with Cartesian coordinates $(X^{\prime},Y^{\prime})$ as
$\begin{bmatrix}X^{\prime}\\ Y^{\prime}\end{bmatrix}=\frac{R}{\sqrt{A(R)}}\sqrt{1-(1-x)^{2}}\begin{bmatrix}\text{cos}(\varphi)\\ \text{sin}(\varphi).\end{bmatrix}$ (51)
Here, $R$ denotes a given radial distance from the origin, while $\varphi$ corresponds to the polar angle in the observer’s plane [^1] For a better understanding of this coordinate representation, the reader is referred to Fig. 7 of [58].. The value of $x$ may be computed via the following approximation
$x$ $=$ $A(R)y\Bigg{\{}1-\frac{e}{100}(1-A(R))y\left[\text{ln}\left(1-\frac{y}{2}\right)+\frac{y}{2}\right]$ (52)
$+\frac{(315I_{1}^{2}-135I_{2}+180A^{2}(R)+120I_{1}+32)}{720}y^{2}$
$\qquad\qquad+\left(\frac{I_{1}}{2}+\frac{1}{6}\right)y\Bigg{\}},$
with $y$ being expressed by the trigonometric relation between the polar angle in the observer’s plane and the observer’s inclination angle $\theta_{o}$ via

### v2

@@ III.2 Constraints to the $\gamma$ parameter from shadow’s radius
at $2\sigma$ deviation. It is worth mentioning that many different geometries have been constrained in the literature following this observation, see e.g. [40, 41, 42, 43, 44] and the review [30].
@@
We note that, to quantify the theoretical biases in the calibration factor utilized in the inference of the shadow’s size of Sgr* A, the EHT collaboration performs simulations of black hole images to study a wide range of effects concerning the accretion flow properties and the underlying space-time. In this regard, the collaboration assumes the Kerr metric by default, although they also cover black hole metrics that either deviate from the Kerr space-time or describe alternative geometries, such as wormholes or boson stars. Because their analysis implicitly assumes asymptotically flat, GR-based solutions, this does not include the Cotton geometry among other theories than stray too much from GR description of gravitational phenomena, by performing tests with the above constraints we introduce a theoretical bias in our own analysis, since we assume these results to be valid as means to set constraints for the CG spacetime. The extent to which the Cotton geometry could modify the above fractional deviation, and the constraints stemming from it, could be evaluated by performing a similar analysis to that by the EHT collaboration. However, such an effort lies well beyond the purpose of this work.
@@
Upon inspection of Fig. 3, we verify that $r_{sh}$ starts to increase around values of $\gamma M\sim 10^{-12}$. Below this value, the CG black hole shadow radius coincides with that of a Schwarzschild black hole. We note that this is qualitatively the same behavior that is evidenced in [30] for the Rindler metric. Actually, the contribution of $\gamma$ to the shadow radius becomes significant when the term $\gamma r\gg 1$, while the upper $2\sigma$ bound in Eq.(24) becomes saturated at a value of approximately $\gamma M\simeq 3.525\times 10^{-12}$. For the values of $\gamma M$ predicted by galaxy radial acceleration rotation curves in [10, 12], or the recent constraints obtained by analyzing CG’s influence on CMB power spectrum and large scale structures [46] (i.e. both estimating $\gamma M\sim 10^{-17}$), the shadow radius of the Cotton black hole lies well within the constraints imposed by Sgr A*’s shadow bounds, though there is no appreciable difference between it and its Schwarzschild counterpart. This result is not surprising, given that $\gamma$ encodes large-distance gravitational effects within the CG framework. Beyond the $\gamma$ parameter range considered in Fig. 3, we also found that the Cotton black hole shadow radius remains within the $2\sigma$ bounds for $\gamma M\sim 10^{20}$. Such values describe extremely large and unphysical gravitational acceleration effects, exceeding any other current estimates by several orders of magnitude. This merited their exclusion from analysis in the following sections.
@@ IV.1 Exact deflection angle
To proceed with our derivation, it is convenient to make the following variable change in the radial coordinate $u=1/r$ in the expression (14). This allows us to re-write the photon equation of motion as
@@
$\alpha=\sqrt{\frac{1}{2M}}\left(\int_{u_{R}}^{u_{0}}\frac{du}{\sqrt{G(u)}}+\int_{u_{S}}^{u_{0}}\frac{du}{\sqrt{G(u)}}\right)+\Psi_{R}-\Psi_{S}.$ (27)
Where $u_{0}=1/r_{0}$ is the inverse of the distance of closest approach, while $u_{R}=1/r_{R}$ and $u_{S}$ are the inverse of the radius of the observer and the source, respectively. The function $\Psi$ encodes a correction term accounting for the finite distance, and is given by
$\Psi=\arcsin\bigl{[}bu\sqrt{A(u)}\bigr{]},$ (28)
where the indices refer to the position of the observer (R) and the source (S). Note that if we take the limit of $\gamma\rightarrow 0$, and consider also that observer and source are located at infinity we will have
@@
which is the well-known deflection angle formula for the Schwarzschild solution. This change in the definition of the deflection angle is necessary because the Cotton metric is not asymptotically flat. Moving on, the integrals of equation (27) can be solved by rewriting the function $G(u)$ as
@@
where $u_{1},u_{2}$ and $u_{3}$ are the roots of the cubic polynomial. Since $u_{0}=1/r_{0}$ is one of the roots of this polynomial, we have chosen (for reasons that will become clear later) $u_{2}=1/r_{0}$. Now we can write
$u^{3}-u^{2}(u_{1}+u_{3}+1/r_{0})+u(u_{1}u_{3}+u_{1}/r_{0}+u_{3}/r_{0})$
$-u_{1}u_{3}/r_{0}=u^{3}-\frac{u^{2}}{2M}-\frac{u\gamma}{2M}+\frac{1}{2Mb^{2}}.$ (31)
@@
$u_{1}+u_{3}+1/r_{0}=\frac{1}{2M},$ (32)
$u_{1}u_{3}+u_{1}/r_{0}+u_{3}/r_{0}=-\frac{\gamma}{2M}.$
Solving the above system we find that $u_{1}$ and $u_{3}$ are given by the following expressions
@@
whose signs cannot be chosen independently. Note that we have two different solutions, one of which is $u_{1}>u_{3}$ and another that $u_{3}>u_{1}$. Since our goal is that these roots fulfill the condition $u_{3}>u_{2}>u_{1}$, we choose the negative sign in the expression $u_{1}$ and the positive sign in the expression $u_{3}$. Now both integrals of (27) can be put into the form
@@
where $F(\delta,k^{2})$ is a incomplete first-order elliptic integral of the parameters
@@
$\Psi(u)=\arcsin\left(bu\sqrt{-2Mu+\frac{\gamma}{u}+1}\right).$ (38)
@@
$\delta(u^{\prime})=\arcsin\Biggl{[}2\sqrt{2}\Biggl{(}-\left(Mr_{0}(u^{\prime}r_{0}-1)\sqrt{r_{0}^{2}c}\right)\times\Biggl{(}\Bigl{(}\sqrt{r_{0}^{2}c}+6Mr_{0}-r_{0}^{2}\Bigr{)}\left(\sqrt{r_{0}^{2}c}-2Mr_{0}(2Rr_{0}+1)+r_{0}^{2}\right)\Biggr{)}^{-1}\Biggr{)}^{1/2}\Biggr{]},$ (39)
@@
$r_{0}=\frac{b^{2}\gamma}{3}+\frac{2}{3}b\sqrt{b^{2}\gamma^{2}+3}\cos\left(\frac{1}{3}\cos^{-1}\left(\frac{\sqrt{-\frac{1}{b^{4}\left(-\gamma^{2}\right)-3b^{2}}}\left(-2b^{6}\gamma^{3}-9b^{4}\gamma+54b^{2}M\right)}{2\left(b^{4}\left(-\gamma^{2}\right)-3b^{2}\right)}\right)\right).$
For the sake of the graphical representation of the deflection angle we prefer to use the variable $b^{\prime}=1-b_{c}/b$. This variable is limited to a value between $0\leq b^{\prime}\leq 1$, where zero represents the strong field limit and one the weak field one. In Figs. 4 and 5 we show the deflection angle as a function of the variable $b^{\prime}$ for values of $\gamma$M in the range from $10^{-7}$ to $1$. We see that $\alpha$ decreases with increasing $\gamma$ and that there is no significant change in the values of $\alpha$ from $\gamma M<10^{-4}$. If we increase the value of the parameter $\gamma M$ beyond the order of magnitude of $10^{-4}$, the deflection angle begins to give negative values. Observationally, this value diverges for $b^{\prime}=0$ and tends to zero for $b^{\prime}=1$. We expect our theoretical result to reproduce this behaviour. In other words, for the phenomenon of light deflection, Cotton’s parameter is limited to values smaller than $\gamma M=10^{-4}$.
Figure 4: Graphical representation of the deflection angle $\alpha$ (in radians) as a function of the reduced impact parameter $b^{\prime}$ for values of the CG parameter $\gamma M=\{10^{-4},10^{-5},10^{-6},10^{-7}\}$.
Figure 5: Graphical representation of the deflection angle $\alpha$ (in radians) as a function of the reduced impact parameter $b^{\prime}$ for values of the CG parameter $\gamma M=\{1,10^{-1},10^{-2},10^{-3}\}$.
@@ V Images of black holes with thin accretion disks
Having restricted the $\gamma$-space parameter, we now proceed to generate the images of the Cotton black hole (by pushing the value saturating the upper $2\sigma$ bound of Sgr A* shadow) and qualitatively compare its appearance to that of a Schwarzschild black hole of the same mass parameter. We shall consider the setting of a black hole illuminated by an optically (i.e. fully transparent to its own radiation) and geometrically thin accretion disk, when observed at both a face-on inclination (i.e. $\theta_{o}=0^{\circ}$) and at a nearly edge-on inclination (i.e. $\theta_{o}=80^{\circ}$).
We begin our considerations from Eq.(14), which we conveniently re-write under the form
$\frac{d\phi}{dr}=\pm\frac{b}{C(r)}\sqrt{\frac{1}{1-b^{2}V_{eff}}},$ (41)
describing the variation in a photon’s azimuthal angle as a function of its radial coordinate. By numerically ray-tracing the previous equation for a bunch of trajectories within some range of the impact parameter $b$, we compute the null geodesics from an observer’s screen backwards towards their point of origin. Fig. 6 displays the geodesics in the Schwarzschild space-time (left panel) for $M=1$, and in the CG space-time (right panel), the latter with a value of $\gamma M=3.525\times 10^{-12}$. Upon inspection of Fig. 6, one highlights the absence of any significant qualitative differences between the strong lensing properties of the two geometries, under the considered parameter values. In addition to larger distances, the effects of $\gamma$ become more noticeable at much higher orders of magnitude than those considered here, as is discussed in section VI below.
Figure 6: Null geodesics in the Schwarzschild space-time (left panel) and in the CG space-time ($\gamma M=3.525\times 10^{-12}$) as a function of impact parameter $b$ normalized with respect to the mass parameter $M$. The photon trajectories are categorized according to the number of half-turns $n$ around the black hole as the ($n=0$) direct image (black), the $n=1$ photon ring (gold), and the $n=2$ photon ring (red), and with an impact parameter spacing of $1/5$, $1/50$, and $1/500$, respectively. The black circle depicts the inner region to the event horizon. The dashed black circumference represents the photon sphere radius. The black hole equatorial plane is represented by the black vertical line. The observer’s screen is located on the right-side axis.
@@
$I(r;\gamma,\mu,\sigma)=\frac{e^{-\frac{1}{2}\left[\beta+\text{arcsinh}(\frac{r-\mu}{\sigma})\right]^{2}}}{\sqrt{(r-\mu)^{2}+\sigma^{2}}},$ (42)
where the $\beta,\mu$ and $\sigma$ parameters modify the skewness, peak location and kurtosis of the distribution, respectively. Such models are designed to reproduce the most prominent image features of time-averaged General Relativistic Magneto-Hydrodynamic (GRMHD) simulations of a thick accretion flow around a Kerr black hole, providing a less computationally taxing alternative towards the generation of black hole images [31]. Given the fact that such parameters are also indirectly dependent on the background geometry (for instance, via the location of its event horizon or photon spheres), they can be also employed to generate simulated images of alternative black holes both within and beyond GR, this way producing templates of expected optical appearances under different assumptions for the accretion flow.
@@
For the purpose of generating black hole images we will restrict out analysis up to the $n=2$ photon ring, as higher-order photon rings will provide a negligible contribution to the net luminosity (as shall be explained in the next section). We shall suitably adapt the three models introduced in [31] to a spherically symmetric scenario, to simulate the accretion flow around the Schwarzschild and Cotton black holes, with the parameters
GLM1 ::: $\beta=-\frac{3}{2}\,,\quad\mu=0\,,\qquad\ \sigma=\frac{M}{2}\ ,$ (44)
GLM2 ::: $\beta=0\,,\qquad\mu=0\,,\qquad\ \sigma=\frac{M}{2}\ ,$ (45)
GLM3 ::: $\beta=-2\,,\quad\ \mu=\frac{17M}{3}\,,\ \sigma=\frac{M}{4}\ .$ (46)
whose corresponding emission profile is depicted in Figure 7.
Figure 7: Emission profiles corresponding to the GLM1, GLM2 and GLM3 models described by the choice of parameters in (44), (45) and (46), respectively, of the intensity function (42).
@@
Figure 8: Visual appearance, in the impact parameter space, of the Schwarzschild black hole ($1^{st}$ and $3^{rd}$ columns) and the CG black hole with ($\gamma M=3.525\times 10^{-12}$) ($2^{nd}$ and $4^{th}$ columns), for the GLM3 (top), GLM1 (middle) and GLM2 (bottom) emission profiles, for an observer located at a distance of $r_{o}\simeq 4.2\times 10^{10}M$ [38], seen from a face-on orientation and from an inclination of $80^{\circ}$, respectively.
@@ VI Isoradials
In this section, we perform an analysis of isoradial curves in the Cotton and Schwarzschild space-times, to complement the black hole images presented in the previous section. These curves refer to the contours, in the observer’s image plane, that lie at a radial constant distance from the black hole’s center.
We solve the isoradial equatorial curves corresponding to the $n=0$ emission, following the analytical approximation provided by [50], which states that the arrival point of a light ray may be mapped onto an observer’s plane with Cartesian coordinates $(X^{\prime},Y^{\prime})$ as
$\begin{bmatrix}X^{\prime}\\ Y^{\prime}\end{bmatrix}=\frac{R}{\sqrt{A(R)}}\sqrt{1-(1-x)^{2}}\begin{bmatrix}\text{cos}\varphi\\ \text{sin}\varphi\end{bmatrix}$ (47)
Here, $R$ denotes a given radial distance from the origin, $x=1-\cos\alpha$, where $\alpha$ is the emission angle, while $\varphi$ corresponds to the polar angle in the observer’s plane [^2] For a better understanding of this coordinate representation, the reader is referred to Fig. 7 of [50].. The value of $x$ may be computed via a refinement of Poutanen approximation developed in [51] via the following formula [50]
$x=A(R)y\left\{1+\left(\frac{I_{1}}{2}+\frac{1}{6}\right)y+\frac{(315I_{1}^{2}-135I_{2}+180A^{2}(R)+120I_{1}+32)}{720}y^{2}-\frac{e}{100}(1-A(R))y\left[\text{ln}\left(1-\frac{y}{2}\right)+\frac{y}{2}\right]\right\},$
with $y=1-\cos\psi$ (with $\psi$ the azimuthal angle) being expressed by the trigonometric relation between the polar angle in the observer’s plane and the observer’s inclination angle $\theta_{o}$ via